# Complex magnetic landscape in kagome metal TbTi3Bi4 and competition with family of Ln2-xTi6+xBi9 (Ln: Tb–Lu) shurikagome metals

### v1

@@ III.2 Properties of TbTi3Bi4
Overall, our work into TbTi3Bi4 has revealed a unique antiferromagnetic, cleavable metal with an exceedingly complex landscape of magnetic phases. The magnetic and transport properties are rich and diverse, with multiple unique magnetic phases to investigate, strong switching from positive to negative magnetoresistance near the 2/3 magnetization plateau, and an unusual coupling between the magnetic phase pockets and the transport. We now turn to briefly investigate the magnetic properties of the Ln2-xTi6+xBi9 shurikagome metals and the influence of disorder on the Ln–Ln chains.
@@
Tb2-xTi6+xBi9 | 0.89 | 0.76 | 1.65 | 1.4 | 7.4 | 6.3 | 7.0 | 11.5 | 9.7 | -24 |
Dy2-xTi6+xBi9 | 0.84 | 0.62 | 1.46 | 1.3 | 6.8 | 5.9 | 6.8 | 7.8 | 9.6 | -14 |
Ho2-xTi6+xBi9 | 0.83 | 0.56 | 1.39 | 1.2 | 6.9 | 6.0 | 6.9 | 3.2 | 9.9 | -0.2 |
Er2-xTi6+xBi9 | 0.81 | 0.54 | 1.35 | 1.2 | 6.1 | 5.3 | 6.2 | 2.4 | 10.6 | -0.9 |
Tm2-xTi6+xBi9 | 0.81 | 0.56 | 1.37 | 1.2 | 4.8 | 4.3 | 5.3 | 1.6 | 7.3 | 14 |
Lu2-xTi6+xBi9 | 0.75 | 0.46 | 1.21 | 1.0 | 0 | 0 | 0.1 | N/A | N/A | N/A |
III.3 Properties of Ln2-xTi6+xBi9
Diffraction studies on the Ln2-xTi6+xBi9 compounds revealed several interesting observations. We first attempted to solve the structure using fully occupied Ln sites (forming fully occupied chains). While R values were actually quite favorable, there existed large amounts of excess electron density around the Ln04 and Ln05 sites. The residual Q-peaks can be eliminated by allowing the occupancy of the Ln sites to vary, which results in partially occupied Ln chains. However, compositions extracted from single-crystal X-ray diffraction (SCXRD) using the assumption of pure (although partially occupied) Ln sites are in strong disagreement with electron dispersive spectroscopy (EDS) results. Bluntly, there is not enough Ti in the unit cell derived from SCXRD. This can be remedied by allowing antisite disorder on the Ln sites, allowing the composition of the Ln to vary, but restraining it such that the site remains fully occupied (e.g. Ln04 + Ti04 = 1). Even this appears to slightly underestimate the amount of Ti, and EDS results suggest that some Ti may be substituting elsewhere as well. This is perhaps expected, as the published Ti–Bi phase diagram indicates a full solid solution over the range of 50:50 Ti:Bi to 35:65 Ti:Bi.[41]
However, the SCXRD data did not independently suggest any additional site substitutions, and the constraints mentioned above yielded quite favorable statistics. Figure 6(a) highlights the cell volume for the Ln2-xTi6+xBi9 shurikagome metals as a function of the 8-coordinate Shannon radius. Despite the variable rare-earth occupancy, the compounds still adhere to a nearly linear expansion in cell volume, as one would expect of a rare-earth series. These results are briefly summarized in Table 1. While CIF files are provided in the ESI, here we focus on the influence of the Ln atom and the disorder in crystals of Ln2-xTi6+xBi9. The columns labeled Occ and Occ indicate the refined occupancy of the chain sites Ln04 and Ln05. As mentioned before, the Ti occupancy on those sites is constrained such that they add to 1. The sum of the SCXRD-determined occupancies is provided in the column labeled “Tot Ln.” This can be directly compared with the composition extracted from EDS measurements in the next column.
SCXRD appears to consistently calculate higher concentrations of Ln as compared to EDS measurements, over approximating by about 15%. Some interesting trends appear in the overall data. Both the Ln04 and Ln05 site occupancies decrease as we move to smaller Ln atoms. This effect is substantially more pronounced on the Ln05 (smaller) rare-earth site, where the occupancy is nearly halved by the time we reach Lu2-xTi6+xBi9. We performed some exploratory syntheses around Tb2-xTi6+xBi9 in an attempt to fully saturate the Ln site, though only small changes in the rare-earth composition were noted.
Considering the high disorder in crystals of Ln2-xTi6+xBi9, particularly within the Ln sublattices, it is not clear what magnetic response to expect. Naively, the most likely scenario was freezing into a spin-glass, or a complete loss of a well-defined magnetic ground state in the non-Kramers ions. We now turn to a brief overview of the magnetic properties of the Ln2-xTi6+xBi9 compounds.
The compounds are tetragonal (P4/nmm) yielding two directions of interest, (parallel to chain) and (perpendicular to chain). We did not consider rotations with in the ab plane for this work. All compounds exhibit weak anisotropy with the exception of Tm. When , Tm2-xTi6+xBi9 exhibits an order of magnitude stronger response. For Tb, Dy, and Ho, data collected with exhibits a cusp in the magnetization, strangely resembling an antiferromagnetic transition. For Er and Tm, the data with exhibits a broad peak instead. All other combinations of elements and directions (e.g. Tb with are featureless, although the magnitude of the magnetization remains comparable to the other direction). These results are summarized in Figure 6(a) which provides an overview of the magnetization for the Ln2-xTi6+xBi9 shurikagome metals along their “favored” direction. Please note that the magnetization results are shown in arbitrary units for ease of comparison. Full data sets are provided in the ESI.
@@ III.3 Properties of Ln2-xTi6+xBi9
Note that the data presented in Figure 6(a) includes both zero-field-cooled (ZFC) and field-cooled (FC) data. The lack of ZFC/FC splitting is surprising, as we may reasonably expect the strong disorder to result in a glassy magnetic state. Furthermore, AC susceptibility measurements on Tb2-xTi6+xBi9 indicate no frequency dependence. However, this is not universal throughout the series, as Tm2-xTi6+xBi9 exhibits a strong frequency response. It is striking that the features follow the temperature scale expected of the latter half of the rare-earth row – Tb possesses the highest ordering temperature, and Tm possesses the lowest.
@@
Figure 6(b) demonstrates the isothermal magnetization for all magnetic Ln2-xTi6+xBi9 compounds at 1.8 K. Again we have opted to show only one direction (the same direction as in Figure 6(a)) for simplicity. All compounds near saturation by 7 T, though Er2-xTi6+xBi9 has not saturated, owing to an unusual inflection in the data near 4 T. Regardless, we can extract the approximate saturation magnetization for all compounds, which has been summarized in Table 1 (column M). Our quantitative analysis assumes the molar susceptibility can be expressed in terms of the ideal formula unit Ln2-xTi6+xBi9. This will underestimate the true saturation magnetization, as we will be normalizing to 2 rare-earth atoms per formula unit instead of (2-x). However, we can compare these results with the expected saturation magnetization from XRD and EDS results (M and M). Comparing directly, we find that the experimental results actually agree quite well with the SCXRD results. Some other parameters of interest (e.g. and ) have also been added to Table 1. Due to the disorder and variable number of Ln sites, we consider these values as first approximations.
Despite the high levels of disorder, some samples of Ln2-xTi6+xBi9 (e.g. Tb) appear to show magnetization consistent with long-range order. This is highly unexpected, and considering the lack of ZFC/FC splitting and no clear frequency dependence, more detailed investigations into the magnetism and chemical disorder of the Ln2-xTi6+xBi9 family are clearly needed.
@@ IV Conclusions
We began this work by describing a modified synthesis procedure that is capable of subverting phase competition to realize TbTi3Bi4, a kagome metal belonging to the AM3X4 family of materials. We described the rational for the formation of TbTi3Bi4, including the increased competition with another class of materials, the Ln2-xTi6+xBi9 shurikagome metals for small rare-earth species (Ln: Tb–Lu). Our work into both classes of materials have demonstrated that TbTi3Bi4 is an exceedingly interesting candidate material – an exfoliable, antiferromagnetic kagome metal with a complex landscape of magnetic phases surrounding multiple metamagnetic transitions. Transport measurements highlight the impact the rare-earth magnetism has on the kagome-dominated transport. Upon entering the magnetization plateau, the system sharply switches between positive (30-40%) and negative (-55%) magnetoresistance in a system that is already quite metallic (1 Ohm-cm). We have also briefly examined the properties of the Ln2-xTi6+xBi9 shurikagome metals. Despite considerable disorder on the rare-earth sublattices, some systems exhibit magnetization responses more indicative of long-range order. Together, these results not only provide insight into a new kagome lattice and an unusual family of disordered magnets, but also illuminate a synthetic strategy to help realize new members of the AM3X4 family.
@@ V Acknowledgements
The work of B.R.O (synthesis, data collection, writing), H.Z. (analysis), K.G. (He3 data collection), A.F.M. (He3 data collection), and M.A.M. (editing), was supported by the U.S. Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES), Materials Sciences and Engineering Division. We thank the X-ray laboratory of the Oak Ridge National Laboratory Spallation Neutron Source for use of their Photonics Laue Camera System.

### v2

@@ III.2 Properties of TbTi3Bi4
Overall, our work into TbTi3Bi4 has revealed a unique antiferromagnetic, cleavable metal with an exceedingly complex landscape of magnetic phases. The magnetic and transport properties are rich and diverse, with multiple unique magnetic phases to investigate, strong switching from positive to negative MR near the 2/3 M plateau, and a complex coupling between the magnetic phase pockets and the transport. We now turn to briefly investigate the magnetic properties of the Ln2-xTi6+xBi9 shurikagome metals and the influence of disorder on the Ln–Ln chains.
III.3 Properties of Ln2-xTi6+xBi9 (Ln: Tb–Lu) family
When we first identified the Ln2-xTi6+xBi9 shurikagome metals, we were surprised by the substitution of Ln into the Ti8Bi9 structure. Unlike integration of the rare-earth elements into frameworks like CoSn, there are no clear interstitial voids in Ti8Bi9. Further, the miscibility of most rare-earth elements into Ti is low, and it is initially unintuitive that Ln is the predominant defect. The incomplete substitution of Ln is also perplexing.
While single crystal X-ray diffraction (SCXRD) results can be solved using fully occupied Ln sites (forming fully occupied chains) with relatively favorable statistics, large amounts of excess electron density accumulate on the Ln4 and Ln5 sites. The residual Q-peaks can be eliminated by allowing the occupancy of the Ln sites to vary, which results in partially occupied Ln chains. This is also consistent with energy dispersive spectroscopy (EDS) results, which support a Ln content from 6–9% (ideal Ln2Ti6Bi9 is 12% Ln). For simplicity, we have restrained the chain sites to remain fully occupied (e.g. Occ + Occ = 1). Even this appears to slightly underestimate the amount of Ti, and EDS results suggest that some Ti may be substituting elsewhere as well. The published Ti–Bi phase diagram supports Ti substitution, as Ti8Bi9 supports a full solid solution from 50:50 Ti:Bi to 35:65 Ti:Bi.[51] Regardless, the SCXRD data did not improve substantially with more complex substitution models. SCXRD tables, CIF files, and EDS results for the entire series can be found in the ESI.
The substitution of the Ln atoms creates a clear separation between the Ln4/Ln5 sites and the Ti-based shurikagome network. As suggested by our designation “shurikagome,” the coordination of the Ti sublattice in the Ln2-xTi6+xBi9 compounds contains pieces of the corner-sharing triangular motif observed in prototypical kagome samples. The parent structure Ti8Bi9 has been known for some time,[50] though the electronic structure has only been calculated through high-throughput compendiums like Materials Project.[59] In an effort to understand the underlying electronic structure of the Ln2-xTi6+xBi9 compounds, we have performed first-principle density functional theory calculations on the ideal Ln2Ti6Bi9 structure. To avoid the additional complication of magnetism, we examine the electronic structure of Lu2Ti6Bi9 as a proxy for the material family.
Figure 6 provides the calculated electronic structure for Lu2Ti6Bi9 alongside a schematic diagram of the P4/nmm Brilluoin zone.[60] There are several interesting features apparent within the data. We first note a Fermi-level band edge and associated “flat–band,” in two distinct directions, centered at (). The extent of this flat band is relatively limited when compared to the prototypical kagome band structure, though the proximity to the Fermi level is interesting. Similar to the case of the LnTi3Bi4 compounds, the states near the Fermi level are almost entirely comprised of Ti-based orbitals arising from shurikagome motif. Beyond the flat-bands, the Lu2Ti6Bi9 band structure contains many dispersive Dirac-like features and several avoided crossings near E (at and near the -point, 10 meV above E between and , 50 meV above E between and ).
Flat bands and/or van Hove singularities near E are commonly cited as the cause for a wide array of unusual behavior, presumably due to the localized nature of electrons in flat-bands and associated correlative effects. The numerous band edges at or very near (25 meV) the Fermi level in Lu2Ti6Bi9 are suggestive of numerous Fermi surface Lifshitz transitions, which supports the off-stoichiometric behavior of these compounds. This additionally indicates a complex, highly sample-dependent Fermi-surface structure. Further, while the spectral weight associated with the Fermi-level “flat-bands” in Lu2Ti6Bi9 is relatively small, it is still suggestive of potential novel behavior potentially observable via ARPES.
Out of curiosity, we performed preliminary ARPES experiments on an in-house laser ARPES system on crystals of Tb2Ti6Bi9 (which has the highest rare-earth occupancy) at approximately 10 K. We were able to observe a high density of states near the Fermi level, consistent with Figure 6 and the associated “flat-bands.” However, the intrinsic disorder in the Ln2-xTi6+xBi9 compounds is currently prohibitive to performing in-depth ARPES investigations, particularly considering the relatively complex electronic structure. Whether the flat bands can be directly ascribed to the Ti-based “shurikagome” sublattice (in analogy to tight binding calculations in the prototypical kagome metals) is a subject of current investigation. We believe that the applicability of both the DFT and ARPES models will require parallel efforts to synthesize compounds with reduced disorder. The existence of the halfnium-based Hf8Bi9[50] and some similar vanadium-based V8Sb9[61] compounds may offer some flexibility in the choice of framework cations that may offer routes to achieve fully-occupied Ln2M6X9 samples.
Indeed, synthesizing fully-occupied Ti8Bi9-based shurikagome compounds may be challenging, as the extent of disorder in samples of Ln2-xTi6+xBi9 appear to vary quite strongly with the rare-earth species. Figure 7(a) highlights the cell volume for the Ln2-xTi6+xBi9 shurikagome metals as a function of the 8-coordinate Shannon radius. Despite the variable rare-earth occupancy, the compounds still adhere to a nearly linear expansion in cell volume, as one would expect of a rare-earth series. These results are briefly summarized in Table 1. The columns labeled Occ and Occ indicate the refined occupancy of the chain sites Ln4 and Ln5. The sum of the SCXRD-determined occupancies is provided in the column labeled “Tot Ln.” This can be directly compared with the composition extracted from EDS measurements in the next column.
Some interesting trends appear in the agglomerate data. First, SCXRD appears to consistently calculate higher concentrations of Ln when compared to EDS. However, regardless of technique, both the Ln4 and Ln5 site occupancies decrease as we move to smaller Ln atoms. This effect is substantially more pronounced on the Ln5 (smaller) rare-earth site, where the occupancy is nearly halved by the time we reach Lu2-xTi6+xBi9. We performed some exploratory syntheses around Tb2-xTi6+xBi9 in an attempt to fully saturate the Ln site, though only small changes in the rare-earth composition were noted. Coincidentally, we have not observed the formation of Ln2-xTi6+xBi9 with large rare-earths (La–Gd). We speculate that the larger rare-earths may destabilize the structure. However, under this assumption, we find it unusual that smaller rare-earth elements have decreasing occupancies, particularly on the smaller site.
Remarkably, despite the high degree of chemical disorder in crystals of Ln2-xTi6+xBi9, magnetization measurements exhibit properties more indicative of antiferromagnetic order. Due to the tetragonal (P4/nmm) structure and the quasi-1D Ln–Ln chain, there are two directions of interest for magnetization measurements, (parallel to chain) and (perpendicular to chain). Samples containing Tb, Dy, and Ho exhibit a cusp in the magnetization when , whereas Er and Tm have features when instead. All compounds exhibit weak anisotropy with the exception of Tm. When , Tm2-xTi6+xBi9 exhibits an order of magnitude stronger response. These results are summarized in Figure 7(b) which provides an overview of the magnetization for the Ln2-xTi6+xBi9 shurikagome metals along their “favored” direction. Note that Figure 7(b) is shown in arbitrary units for ease of comparison, though full data sets are provided in the ESI.
@@ III.3 Properties of Ln2-xTi6+xBi9 (Ln: Tb–Lu) family
Tb2-xTi6+xBi9 | 0.89 | 0.76 | 1.65 | 1.4 | 7.4 | 6.3 | 7.0 | 11.5 | 9.8 | 21 |
Dy2-xTi6+xBi9 | 0.84 | 0.62 | 1.46 | 1.3 | 6.8 | 5.9 | 6.8 | 7.8 | 9.6 | 14 |
Ho2-xTi6+xBi9 | 0.83 | 0.56 | 1.39 | 1.2 | 6.9 | 6.0 | 6.9 | 3.2 | 9.9 | 0.8 |
Er2-xTi6+xBi9 | 0.81 | 0.54 | 1.35 | 1.2 | 6.1 | 5.3 | 6.2 | 2.4 | 10.5 | 1.0 |
Tm2-xTi6+xBi9 | 0.81 | 0.56 | 1.37 | 1.2 | 4.8 | 4.3 | 5.3 | 1.6 | 8.6 | 1.1 |
Lu2-xTi6+xBi9 | 0.75 | 0.46 | 1.21 | 1.0 | N/A | N/A | N/A | N/A | N/A | N/A |
@@
The data presented in Figure 7(b) includes both ZFC and FC data for Tb–Er samples. The lack of ZFC/FC splitting is surprising, as we may reasonably expect the strong disorder to result in a glassy magnetic state. Furthermore, AC susceptibility measurements on Tb2-xTi6+xBi9 indicate no frequency dependence. However, this is not universal throughout the series, as Tm2-xTi6+xBi9 exhibits some frequency dependence. Notably, however, Tm also possesses one of the lowest Ln occupancies (Tm1.4Ti2.6Bi9) across the series. The magnetization anomalies follow the temperature scale (de Gennes scaling) expected of the latter half of the rare-earth row, with Tb possessing the highest ordering temperature, and Tm possessing the lowest. The results are generally reminiscent of antiferromagnetic-like order, particularly considering the lack of ZFC/FC splitting. As an additional diagnostic tool, we performed zero-field heat capacity measurements (see ESI) on both Tb2-xTi6+xBi9 and Er2-xTi6+xBi9 crystals from 300 K to 1.8 K. Both samples show clear, albeit small, heat capacity anomalies. These results are in good agreement with magnetization results and further support some degree of non-glassy magnetic order in the series.
@@
Figure 7(c) demonstrates the isothermal magnetization for all magnetic Ln2-xTi6+xBi9 compounds at 1.8 K. For simplicity we have opted to show the same field directions as in Figure 7(b), though both the and directions are included in the ESI. All compounds near saturation by 7 T, though Er2-xTi6+xBi9 has not saturated, owing to an unusual inflection in the data near 4 T. Regardless, we can extract the approximate saturation magnetization for all compounds, which has been summarized in Table 1 (column M). Our quantitative analysis assumes the molar susceptibility can be expressed in terms of the ideal formula unit Ln2Ti6Bi9. This underestimates the true saturation magnetization, though it can be compared directly with the “expected saturation magnetization from XRD and EDS results (M and M). Some other parameters of interest (e.g. and ) have also been added to Table 1. Of course, due to the disorder and variable number of Ln sites, we consider these values as first approximations.
Remarkably, despite the high levels of disorder, the majority of Ln2-xTi6+xBi9 appear to show magnetic properties consistent with non-glassy magnetic order. The lack of ZFC/FC splitting, the lack of frequency dependence in most samples, and the presence of well-defined heat capacity anomalies suggest some degree of conventional magnetic ordering. The disorder assuredly plays a non-trivial role in the manifestation of this order, and we cannot completely exclude the possibility of some glassiness. More detailed investigations into the magnetism and chemical disorder of the Ln2-xTi6+xBi9 family are clearly needed. Future work should focus on controlling chemical disorder in the shurikagome metals, particularly with an emphasis on realizing an idealized Ln2M6X9 analog.
@@ IV Conclusions
We began this work by describing a modified synthesis procedure that is capable of subverting phase competition to realize TbTi3Bi4, the smallest rare-earth containing member of the LnTi3Bi4 kagome metals. We described the rationale for the formation of TbTi3Bi4, including the increased competition with the shurikagome metals Ln2-xTi6+xBi9 (Ln: Tb–Lu). Our work into both classes of materials have demonstrated that TbTi3Bi4 is an exceedingly interesting candidate material – an exfoliable, antiferromagnetic kagome metal with a complex landscape of magnetic phases, metamagnetic transitions, and 1/3 and 2/3 magnetization plateaus. Transport measurements highlight the impact the rare-earth magnetism has on the kagome-dominated transport. Upon entering the magnetization plateau, the system sharply switches between positive (+40%) and negative (50%) magnetoresistance in a system that is already quite metallic (1 Ohm cm). We have also briefly examined the properties of the Ln2-xTi6+xBi9 shurikagome metals. Despite considerable disorder on the rare-earth sublattices, the systems exhibit magnetization responses reminiscent of non-glassy antiferromagnetic order. Potential routes towards fully-occupied Ln2M6X9 analogs may provide a cleavable system with an interesting electronic structure. Together, our results not only provide insight into a new kagome lattice and an unusual family of chemically-disordered magnets, but also illuminate a synthetic strategy to help realize new members of the AM3X4 family.
@@ V Acknowledgements
This work was supported by the U.S. Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES), Materials Sciences and Engineering Division. We thank Jong Keum and the X-ray laboratory of the Oak Ridge National Laboratory Spallation Neutron Source (SNS) for use of their Photonic Science Laue camera. Research directed by G.D.S. is additionally sponsored by the Laboratory Directed Research and Development Program of Oak Ridge National Laboratory, managed by UT-Battelle, LLC, for the US Department of Energy. This research used resources of the Compute and Data Environment for Science (CADES) at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725. Portions of this work were supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Science Center (ARPES, Q.L., R.G.M.)
VI Supplementary Info.
Supplementary crystallography tables and corresponding CIF files; Additional magnetic phase diagram utilizing zero-field-cooled virgin curves; Preliminary ARPES data on TbTi3Bi4; Additional magnetoresistance data with full temperature-field data set and full field loops; Table of shurikagome compositions measured by EDS; Preliminary ARPES data on Tb2-xTi6+xBi9; Magnetization versus temperature data for Ln2-xTi6+xBi9 to scale; Isothermal magnetization data for Ln2-xTi6+xBi9; Curie-Weiss fits for Ln2-xTi6+xBi9; Select heat capacity results for Tb2-xTi6+xBi9 and Tm2-xTi6+xBi9